Nick walked a total of 30 blocks in 5 days. Each day he walked 3 blocks more than the previous day. How many blocks did he walk on Day 1.
I think the answer is he did not walk any distance on day 1. Please explain the solution.

Answer by RAY100(1637)

Let x= distance for day 1,,,subsequent days are,,x+3,,x+6,,x+9,,,x+12
.
as total for 5 days is 30,,
.
x +x+3 +x+6 +x+9 =x+12 =30
.
5x +30 =30
.
5x =0
.
x= 0,,,,,,,or he doesn't have to walk at all 1st day,,
.
.
check,,0+3+6+9+12 = 30,,,,ok
.